OpenStudy (anonymous):

Find the solutions to the quadratic equation using the quadratic formula: x^2 - x = 12 Write the answers as integers.

jhonyy9 (jhonyy9):

factoriz it like (x-4)(x+3)

jhonyy9 (jhonyy9):

so than you like using discriminant hope you know sormula of ...

jhonyy9 (jhonyy9):

formula of sorry

OpenStudy (anonymous):

x+x-x-12=0

jhonyy9 (jhonyy9):

so you have learned about discriminant D=b^2 -4ac ?

OpenStudy (anonymous):

kinda.. don't understand it..

jhonyy9 (jhonyy9):

this mean that you dont learned again ?

OpenStudy (anonymous):

I do the whole. x+x-x-12=0 think

OpenStudy (anonymous):

x^2-x-12=0 ... quadratic equation D = b2 - 4ac = 49 D>0 ... The equation has two distinct real roots x1 = 4 x2 = -3 P = {4; -3} http://www.hackmath.net/en/calculator/quadratic-equation?a=1&b=-1&c=-12

OpenStudy (anonymous):

YES! That! Thank you!